OpenStudy (anonymous):

help Solve for x: −2(x + 3) = −2(x + 1) − 4.

Vocaloid (vocaloid):

start by distributing

OpenStudy (anonymous):

I tried I just do not get it

OpenStudy (anonymous):

Distribute.( -2x -6) = (-2x -2) -4 Simplify. x-6= -6 Add 6 to both sides x=12 (not sure how correct this is, its late and i;m tired)

OpenStudy (anonymous):

If you added 6 to both sides it will be x=0